mathslover (mathslover):

A deer wants to save her life from a lion. The lion follows a path whose equation is \(x^2+y^2=16\) . For saving life, the deer moves on the path whose equation is/are ?

mathslover (mathslover):

Please refresh the page friends. I have done some editing.

Parth (parthkohli):

A path that does not intersect with \(x^2 + y^2 = 16\), right?

Parth (parthkohli):

By observation, this is just a circle with radius \(4\) and center \((0,0)\). So we can simply write down the equation of the vertical line \(x = 5\), and voila!

Parth (parthkohli):

\(y = 5\), the horizontal line, works too.

mathslover (mathslover):

\(x^2+y^2 = 16\) would form a circle with radius 4 units, centered at origin.

Parth (parthkohli):

Yup.

mathslover (mathslover):

Yeah! I can also draw a circle with small radius : \(x^2 + y^2 = 4\) ..

Parth (parthkohli):

So the equation \(x = 5\) works, and so does \(y = 5\) :-)

mathslover (mathslover):

or just greater radius \(x^2 + y^2 -64 =0 \) or \(x^2 + y^2 = 64\) Thanks @ParthKohli 

mathslover (mathslover):

well i just imagine, it would be so funny.. a lion moving round a circle and similarly a deer moving round a smaller circle1

Parth (parthkohli):

Indeed.

Parth (parthkohli):

lol

mathslover (mathslover):

Oh! It would be so idiotic ... for the lion... (: . Thanks @ParthKohli .

Parth (parthkohli):

Robotic lions who don't apply their intelligence and use the concept of scattering are rare. :-P

mathslover (mathslover):

:-P mainly found in Gir... ;)

Parth (parthkohli):

lol, do you have any more questions today? I have to go to sleep early today, but I'd delay mine because of your questions. They are pretty interesting :-P

mathslover (mathslover):

:-P mechanics questions are always interesting. Well, I have but I am doing them currently. You go to sleep better. It would take time. Have a great day!

mathslover (mathslover):

^ night!